Circular No. 8118

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2000 QD_{181} = 1986 R1 = 1993 WU (RUSSELL-LINEAR)

R. H. McNaught, Siding Spring Observatory, reports observations of a comet found in Sept. 1986 by K. S. Russell on a 90-min exposure taken by F. G. Watson earlier that month with the U.K. Schmidt Telescope. Unsuccessful attempts were made by Russell, and later by McNaught, to locate the comet on the 30-min follow-up exposure by M. Hartley obtained on Sept. 25. T. B. Spahr, Minor Planet Center, has recently identified the comet with 2000 QD_{181}, an apparently asteroidal object observed by LINEAR on 2000 Aug. 31 and Sept. 5 (cf. MPS 18353), and itself linked by Spahr (MPO 9348) to another LINEAR discovery 2000 XV_{43} (observations Nov. 2000–Jan. 2001 on MPS 23109 and 25364), as well as to 1993 WU, recorded by C. S. Shoemaker et al. with the 0.46-m Palomar Schmidt telescope on 1993 Nov. 19 and 20 (MPS 397), the appearance again being evidently asteroidal. With the knowledge of the clearly cometary orbit, McNaught and M. A. Read have now located and measured the object on the Hartley follow-up plate. The 1986 observations, which have been given the designation P/1986 R1, and orbital elements for the respective epochs 1986 Nov. 26.0, 1993 Oct. 20.0, and 2000 Aug. 4.0 TT, follow:

1986 UT α_{2000} δ_{2000} m_{1} Observer
Sept. 3.64300 0^{h}47^{m}48.^{s}78 −32^{o}23′08′′.3 17 Watson
3.70550 0 47 46.61 −32 23 13.8  _{”}
25.62655 0 27 27.46 −31 06 55.8  Hartley
25.64738 0 27 25.95 −31 06 43.5  _{”}
T = 1986 Dec. 4.8062 TT  ω = 357.^{o}7270
e = 0.557259  Ω = 39.0554 }2000.0
q = 1.595504 AU  i = 20.7524
a = 3.603700 AU n^{o} = 0.1440725 P = 6.841 years
T = 1993 Oct. 11.5441 TT  ω = 357.^{o}7477
e = 0.555720  Ω = 39.0418 }2000.0
q = 1.604777 AU  i = 20.7196
a = 3.612080 AU n^{o} = 0.1435714 P = 6.865 years
T = 2000 Aug. 17.6199 TT  ω = 357.^{o}7420
e = 0.557054  Ω = 39.0471 }2000.0
q = 1.596837 AU  i = 20.7496
a = 3.605034 AU n^{o} = 0.1439926 P = 6.845 years
2003 April 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green